OpenStudy (anonymous):

Which pair of equations represents two perpendicular lines? (Points : 3) Option A: -3x +2y = 10 and 3y = 2x +12 Option B: 2x + 5y = 45 and y + 2/5x = -9 Option C: x= 4y +4 and x +4y=4 Option D: 7x + 4y= 20 and y-3=4/7x

OpenStudy (campbell_st):

if you know the slope of 1 line slope = m then the perpendicular line has a slope if \[slope_{perpendicular} = - \frac{1}{m}\] so for each option, find the slope of the 1st line, then get the negative reciprocal and see if its the slope of the 2nd line Option A -3y + 2y = 10 or y = 3/2 x + 5 slope = 3/2 negative reciprocal is -2/3 2nd line y = 2/3x + 4 slope is 2/3 so not perpendicular... hope it makes sense

OpenStudy (anonymous):

it makes sense im just not sure how to solve to find the slope

OpenStudy (campbell_st):

so you need the line in slope intercept form y = mx + b so just make y the subject in each line option B has y = -2/5x + 9 and y = -2/5x - 9 same value